OpenStudy (anonymous):

If 6 is 24% of a number, what is 40% of the same number? (A) 8 (B) 10 (C) 15 (D) 20

OpenStudy (kropot72):

One percent of the number is 6/24. 40% of the number is (6/24) * 40 = ?

OpenStudy (anonymous):

(6/24) * 40 solve this

OpenStudy (calculusfunctions):

Could you people please stop giving out answers and help.

OpenStudy (calculusfunctions):

Anyone can give out answers but only those are truly blessed with communicating valuable knowledge are true teachers.

OpenStudy (anonymous):

How does an answer require you say solve this.

OpenStudy (anonymous):

Just saying calculaus

OpenStudy (anonymous):

6/24 * 40 ....6 * 40 = 240 and 240/24 is 10 ..so your ANSWER is B 10. I am not a teacher. And I don't plan to be.

OpenStudy (calculusfunctions):

Have you seen any of my lessons on this site?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

That is revelant..how?

OpenStudy (theviper):

Let the no. be \(x.\) Then its 24% is 6 that is \[\frac{24}{100}\times x=6\] \[x=\frac{6\times100}{24}=\frac{600}{24}=25\] Therefore 25 is the no. & its 40% is \[\frac{40}{100}\times25=\frac{1000}{100}=10\] Therefore \(\color{red}B\) is the right answer :)

OpenStudy (anonymous):

OH NO..someone gave an answer....arrest them calc.

OpenStudy (calculusfunctions):

Again just giving away the answers. If you stop giving out answers and allow me to teach perhaps you may see the relevance of my pont.

OpenStudy (anonymous):

Best answer comes from TheViper ..side note (Its not best "lesson on how to do it, you figure it out even though "most lol" people spend 12+ years learning lessons not one post on a website"

OpenStudy (calculusfunctions):

Now you're simply being facetious.

OpenStudy (anonymous):

No...really?